I have been given a question with a given a expected probability of 0.45 for finding a job out of 5 possible trials, how do I find the mean and standard deviation
1
Expert's answer
2020-04-09T14:06:23-0400

This is a binomial distribution case because there are 2 events. Finding a job with probility p and not finding a job with probability q=1-p.

n are the number of trials

Mean = np

=5*0.45

=2.25

Standard deviation ="\sqrt{npq}"

= "\sqrt{5*0.45*(1-0.45)}"

=1.1124